What decimal value will be in eax after this code executes?

   sub eax, eax

   sub ebx, ebx

looptop: cmp ebx, 5

   jge exitloop

   add eax, ebx

   inc ebx

   jmp looptop

exitloop:
Answer #1

Instruction

Step1

Step2

Step3

Step4

Step5

Step6

sub eax, eax

eax –eax =0

sub ebx, ebx

ebx –ebx =0

looptop: cmp ebx, 5

Comparing ebx value with 5

Comparing ebx value with 5

Comparing ebx value with 5

Comparing ebx value with 5

Comparing ebx value with 5

Comparing ebx value with 5

jge exitloop

If 0 is greater than or equal to 5 then jump to “exitloop”

If 1 is greater than or equal to 5 then jump to “exitloop”

If 2 is greater than or equal to 5 then jump to “exitloop”

If 3 is greater than or equal to 5 then jump to “exitloop”

If 4 is greater than or equal to 5 then jump to “exitloop”

If 5 is greater than or equal to 5 then jump to “exitloop”

jump to “exitloop”

add eax, ebx

eax = eax +ebx eax = 0+0 à 0

eax = eax +ebx

eax = 0+1 à 1

eax = eax +ebx

eax = 1+2 à 3

eax = eax +ebx

eax = 3+3 à 6

eax = eax +ebx

eax = 6+4 à 10

inc ebx

ebx = ebx +1

ebx = 0 +1 à 1

ebx = ebx +1

ebx = 1 +1 à 2

ebx = ebx +1

ebx = 2 +1 à 3

ebx = ebx +1

ebx = 3 +1 à 4

ebx = ebx +1

ebx = 4 +1 à 5

jmp looptop

Jump to looptop

Jump to looptop

Jump to looptop

Jump to looptop

Jump to looptop

exitloop:

exitloop

Decimal value will be in eax after this code executes 10